OpenStudy (mathmath333):

In how many ways can a selection can be made of at least one fruit out of 5 bananas,4 mangoes and 4 almonds ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{In how many ways can a selection can be made of at least}\hspace{.33em}\\~\\ & \normalsize \text{one fruit out of 5 bananas,4 mangoes and 4 almonds ? }\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (steve816):

5 * 4 * 4 = 80 ways

OpenStudy (mathmath333):

answer given is 149

OpenStudy (steve816):

What the hell? Let me think...

OpenStudy (♪chibiterasu):

Keep in mind that an almond is not a fruit.

OpenStudy (steve816):

LOL! Did they really have to trick us by putting almond xD

OpenStudy (amistre64):

since when is an almond not a fruit?

OpenStudy (amistre64):

mathmatikers are not botinists :)

OpenStudy (♪chibiterasu):

The almond is the seed of the almond fruit, the drupe.

OpenStudy (amistre64):

i smell a mole ...

OpenStudy (amistre64):

at least ... i cant make sense of the questions to be honest.

OpenStudy (mathmath333):

is the answer 13C1

OpenStudy (steve816):

it can't be 13C1

OpenStudy (anonymous):

13C1?

OpenStudy (phi):

I am thinking you can choose from 0 to 5 bananas so 6 choices 0 to 4 mangoes so 5 choices 0 to 4 almonds so 5 choices 6*5*5= 150 that includes the choice of 0 bananas, 0 mangoes , and 0 almonds so 149 with "fruits" if almonds are fruits

OpenStudy (phi):

if almonds are not fruits, this gives 150 - 5 (subtract off 0,0,0; 0,0,1; ...; 0,0,4) for 145